OpenStudy (anonymous):

why is this not valid for matrices (A+B)(A+B)=A^2+2AB+B^2

OpenStudy (phi):

does AB = BA for matrices?

OpenStudy (anonymous):

i am not sure does it?

OpenStudy (anonymous):

so i doesnt equal

OpenStudy (phi):

Yes. the question even hints at the problem: why is this not valid for matrices so you know something is wrong. AB≠BA is a good guess. when you multiply out (A+B)(A+B) you get A*A + A*B + B*A +B*B if A and B were numbers then A*B is the same as B*A and you could combine them but matrices don't let you do that

OpenStudy (anonymous):

so how do i show this by picking a random matrix and show that AB not equal to BA

OpenStudy (phi):

try \[\left[\begin{matrix}1 & 2 \\ 1 & 1\end{matrix}\right]\] and \[\left[\begin{matrix}1 & 1 \\ 1 & 0\end{matrix}\right]\]